OpenStudy (anonymous):

How many and of what type are the solutions of a quadratic equation when the value of the radicant is 25? a. No real solutions b. Two identical rational solutions c. Two different rational solutions d. Two irrational solutions

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] The radicand of the quadratic formula is called the Discriminant \(\Delta =b^2-4ac\) if \(\Delta=0\) there is only one solution to the quadratic equation if \(\Delta>0\) there are two solutions if \(\Delta <0\) the solutions are complex

OpenStudy (unklerhaukus):

another way of putting this if \(Δ=0\) there is two equal solutions if \(Δ>0\) there are two different solutions if \(Δ<0\) there are no real solutions

OpenStudy (unklerhaukus):

if the discriminant is a square number say\[\Delta = d^2\] then taking the square root will not leave any irrational component \[x_{1,2}=\frac{-b\pm \sqrt\Delta}{2a}=\frac{-b\pm \sqrt d^2}{2a}=x_{1,2}=\frac{-b\pm d}{2a}\]

OpenStudy (anonymous):

so I still don't get the answer im confused

OpenStudy (unklerhaukus):

ok well you have \[\Delta =25\] right?

OpenStudy (anonymous):

no \[\sqrt{25}\]

OpenStudy (unklerhaukus):

(the terminology is a little confusing) but the radicand is the bit Under the square root

OpenStudy (anonymous):

ok so whats the answer

OpenStudy (unklerhaukus):

first check is the radicand equal to, greater than, or less that, zero ?

OpenStudy (anonymous):

greater?

OpenStudy (unklerhaukus):

good, so the must be two different solutions,

OpenStudy (anonymous):

ok thx

OpenStudy (unklerhaukus):

but are they rational or irrational? is 25 a square number?

OpenStudy (unklerhaukus):

do you remember what a square number is ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5

OpenStudy (unklerhaukus):

right. so the quadratic formula would have simplified , the solutions are \[x_{1,2}=\frac{-b\pm \sqrt{25}}{2a}=\frac{-b\pm \sqrt 5^2}{2a}=\frac{-b\pm 5}{2a}\] which means \(x=\frac{-b+5}{2a}\) or \(x=\frac{-b-5}{2a}\) which are both rational because the radical symbol has gone

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

does it make any sense?

OpenStudy (anonymous):

yes thanks!

OpenStudy (unklerhaukus):

Say if a quadratic equation was \[x^2+5x=0\] \(a\) would be 1 \(b\) would be 5 and \(c\) would be 0 so the discriminant would equal 25 as we have just had we found that there would be two different rational solutions this is consistent with if we had factored \[x^2+5x=0\\ x(x+5)=0\] where we find the roots as \(x=0\) or \(x=-5\)